Problem 1.26 Prove that the divergence of a curl is always zero. Check it for function Va in Prob. 1.15
Answer #1

Let old{A}=A_xhat{x}+A_yhat{y}+A_zhat{z} be a vector function.

Curl of old{A} is abla imesold{A}=egin{vmatrix} hat{x} & hat{y} & hat{z} rac{partial}{partial x} & rac{partial}{partial y} & rac{partial}{partial z} A_x & A_y & A_z end{vmatrix}

Divergence of old{A} is ablacdotold{A}=rac{partial A_x}{partial x}+rac{partial A_y}{partial y}+rac{partial A_z}{partial z}

Э.dz Э.4:

Divergence of curl of old{A} is ablacdot( abla imesold{A})=rac{partial}{partial x}left[ rac{partial A_z}{partial y}-rac{partial A_y}{partial z} ight ]+rac{partial}{partial y}left[ rac{partial A_x}{partial z}-rac{partial A_z}{partial x} ight ]+rac{partial}{partial z}left[ rac{partial A_y}{partial x}-rac{partial A_x}{partial y} ight ]

ablacdot( abla imesold{A})=left[ rac{partial^2 A_z}{partial xpartial y}-rac{partial^2 A_y}{partial xpartial z} ight ]+left[ rac{partial^2 A_x}{partial ypartial z}-rac{partial^2 A_z}{partial ypartial x} ight ]+left[ rac{partial^2 A_y}{partial zpartial x}-rac{partial^2 A_x}{partial zpartial y} ight ]

Since order of partial differentiation can be changed,

ablacdot( abla imesold{A})=left[ rac{partial^2 A_z}{partial xpartial y}-rac{partial^2 A_y}{partial zpartial x} ight ]+left[ rac{partial^2 A_x}{partial ypartial z}-rac{partial^2 A_z}{partial xpartial y} ight ]+left[ rac{partial^2 A_y}{partial zpartial x}-rac{partial^2 A_x}{partial ypartial z} ight ]

ablacdot( abla imesold{A})=left[ rac{partial^2 A_z}{partial xpartial y}-rac{partial^2 A_z}{partial xpartial y} ight ]+left[ rac{partial^2 A_x}{partial ypartial z}-rac{partial^2 A_x}{partial ypartial z} ight ]+left[ rac{partial^2 A_y}{partial zpartial x}-rac{partial^2 A_y}{partial zpartial x} ight ]=0

Hence divergence of a curl is always zero.

old{v}=x^2hat{x}+3xz^2hat{y}-2xzhat{z}

Curl of old{v}=x^2hat{x}+3xz^2hat{y}-2xzhat{z} is

abla imesold{v}=egin{vmatrix} hat{x} & hat{y} & hat{z} rac{partial}{partial x} & rac{partial}{partial y} & rac{partial}{partial z} v_x & v_y & v_z end{vmatrix}=hat{x}left[ rac{partial v_z}{partial y}-rac{partial v_y}{partial z} ight ]+hat{y}left[ rac{partial v_x}{partial z}-rac{partial v_z}{partial x} ight ]+hat{z}left[ rac{partial v_y}{partial x}-rac{partial v_x}{partial y} ight ]

Э(-2.rz) 013.rz?)1.^P(2.2) д(-222 )].. Гд(3r?) д(2.2) ду

abla imesold{v}=hat{x}left[ 0-6xz ight ]+hat{y}left[ 0+2z ight ]+hat{z}left[ 3z^2-0 ight ]=-6xzhat{x}+2zhat{y}+3z^2hat{z}

Divergence of curl of old{v}=x^2hat{x}+3xz^2hat{y}-2xzhat{z} is

(-6.rz) д(29 ду . Э(3:2) ,